Lab Activity for RNA Processing: p53 Isoforms Isoforms, simply defined, are different proteins created from the same genomic three part question!!!!
Answer #1

answers:

a) the human p53 gene expresses not only 1 but 12 different p53 isoforms (proteins) which is due to the alternative initiation of translation, alternate splicing, alternative promotor usage. they are expressed in normal human tissues expression but they are abnormally expressed in a wide range of cancer type.

  1. p63 : expresses at least three alternating spliced c-terminal isoforms
  2. p73 :expresses at least seven alternating spliced c-terminal isoforms

they both share a significant structural and functional similarity with the p53. +

  • p53 , p53\beta ,p53\gamma \Delta 40p53\alpha  \Delta40p53\beta
  • \Delta40p53\gamma  \Delta133p53\alpha
  • \Delta133p53\beta \Delta 133p53\gamma
  • \Delta160p53\alpha \Delta 160p53\beta \Delta 160p53\gamma

ans 2): mechanism by which each of these are made:

  • \Delta133p53\alpha, \Delta 133p53\betaand \Delta 133p53\gamma: transcription of p53 mRNA from the internal promoter (intron 4) directed to alternative splicing of exon 9b.
  • \Delta160p53\alpha, \Delta 160p53\beta and \Delta 160p53\gamma : encoded by \Delta 133p53\alpha, \Delta 133p53\betaand \Delta 133p53\gammamRNAs by the alternative initiation of translation at ATG 160, which lies within a sequence environment matching Kozack's consensus criteria.
  • p53\beta : alternative splicing of intron 9 (exon 9b)
  • p53\gamma: truncation of p53 protein of the last 60 amino acids and containing additional 15 amino acids.
  • \Delta40p53\alpha, \Delta 40p53 \beta and \Delta 40p53 \gamma : obtained by classic splicing of the C-terminus and alternatively spliced c-terminal isoforms.

ans 3)

proteins which are translated from alternatively spliced mRNAs will have difference in their amino acids sequences therefore they will have difference in their functional properties.